OpenStudy (anonymous):

Simplify (7x+2)(x^2-3x+4

OpenStudy (jmark):

(7x+2)(x^2-3x+4) (7x+2) (x^2-4x+1x+4) =0 // factories the term // (7x+2) (x(x-4) – 1(x-4)) =0 (7x+2) (x-4) (x-1) =0 (7x+2) = 0, (x-4) = 0, (x-1) =0 x = -2/7, x = 4, x = 1 Class 6 ICSE syllabus - http://goo.gl/tkxh4g